# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 5.4. Prompting Strategies
With the construction of prompts for recommendation tasks, prompting strategies are leveraged to further elicit the general abilities of LLMs in text comprehending and problem solving (Zhao et al., 2023; Liu et al., 2023f; Le Scao and Rush, 2021). In this section, we first summarize prompting strategies specialized for recommender systems, and provide empirical findings on two research questions.
@@ 5.4.1. Prompting Strategies Specialized for Recommender Systems
In general tasks with LLMs, researchers use methods such as Chain-of-Thought (CoT) prompting to stimulate the logical reasoning ability of LLMs for solving complex problems (Diao et al., 2023; Liu et al., 2022). As for ProLLM4Rec, leveraging LLMs for recommender systems can also employ prompting strategies to further improve the performance of recommendations (Yao et al., 2023; Liu et al., 2023e). In addition, due to the user and item settings of the recommendation task, the prompting strategy needs to be customized based on the specific user needs in recommender systems. In this paper, we concentrate on typical prompting strategies for recommendation tasks, i.e., zero-shot prompting, few-shot prompting, recency-focused prompting, role prompting, chain-of-thought prompting and self-prompting strategy, and more advanced planning strategies in ProLLM4Rec are left for future exploration (Liu et al., 2023f).
@@
$\bullet$ Zero-shot prompting is the basic scenario and common form among various prompting strategies. Zero-shot prompting directly provides the context information and task description for LLMs without reference examples. The task description and prompt formats vary in line with different recommendation tasks (Wang et al., 2023c; Liu et al., 2023c; Dai et al., 2023a).
@@
$\bullet$ Few-shot prompting is opposite to zero-shot prompting, and it provides a few demonstrations in prompts to help LLMs better understand the user intention (Brown et al., 2020). While for recommendation tasks, the matching of provided examples with the current interest of personalized users determines the effectiveness of few-shot prompting. Therefore, it is worth discussing how to provide reliable context for subsequent recommendations (Hou et al., 2023).
@@
$\bullet$ Recency-focused prompting is first proposed by (Hou et al., 2023) based on the fact that the next predicted item in recommendations has a greater correlation with the recently interacted item than other historical items. Therefore, explicitly emphasizing the recently interacted items in prompts is also a practical prompting strategy. In our initial prompt, recency-focused prompting is used by default.
@@
$\bullet$ Role prompting refers to playing a “role-playing” game with the language model. In the recommendation task, we can specify the role of LLMs as the recommender system to serve users in a targeted manner, adding auxiliary information to describe the role of recommenders in detailed prompts such as conversational recommender systems (He et al., 2023; Wang et al., 2023h; Jin et al., 2023).
@@
$\bullet$ Chain-of-thought prompting is also known as the CoT prompting, which is widely applied in reasoning tasks such as question answering and mathematical inference (Wei et al., 2022). CoT prompting can elicit the ability of LLMs to solve problems step by step (Kojima et al., 2022), or further explicitly decompose the reasoning and analysis process of the task using least-to-most prompting strategy (Zhou et al., 2022). In the field of recommender systems, explicit steps can also be provided manually by researchers to assist in solving recommendation tasks (Wang and Lim, 2023). For ProLLM4Rec, we not only focus on the basic CoT prompts, but also explore the role of custom-designed step-by-step prompts.
@@
$\bullet$ Self-prompting strategy means that the knowledge for answering questions can be obtained by prompting LLMs multiple times. LLMs can be required to generate relevant knowledge, providing necessary information for concepts in the original problem (Liu et al., 2022). Meanwhile, the randomness and self consistency (Wang et al., 2022) generated by LLMs allow them to generate multiple inference chains, and then use the majority voting method on the results obtained from all chains as final predictions. When using LLMs to the re-ranking stage of a recommender system, there are many possible combinations of candidate items, and the recommendation results from the same language model with the same prompt can be totally different. Therefore, bootstrapping with multiple trials is a scientific guarantee to reduce bias (Hou et al., 2023; Ma et al., 2023; Li et al., 2023i).
@@ 5.4.2. Research Questions and Experimental Setup
In this section, we explore the construction and design of prompting strategies for ProLLM4Rec, and provide empirical findings on the following two research questions:
  • RQ1: Whether different expressions of prompts lead to similar recommendation performance?
  • RQ2: How to choose the planning module for stimulating the recommendation ability of LLMs?
As for the strategy of prompt engineering in ProLLM4Rec, we conduct experiments with different variants of prompts compared with the basic prompt as follows:
@@
  • w/o CoT (step-by-step): we remove the magical spell “let’s think step by step” in the original prompt.
@@
  • w/ ICL (self): we use the last interaction of the same user as the example for In-Context Learning (ICL) (Hou et al., 2023).
  • w/ ICL (others): we retrieve an example from historical sequences of other users for the current user, and add the demonstration to the original prompt. For the selection of examples, we encode recent items of users by textual representations, and search for similar user sequences based on the inner product between vectors.
Table 12. Performance comparison on the prompt engineering of LLMs in ProLLM4Rec.
MovieLens-1M Amazon-Books
Method ndcg@1 ndcg@10 ndcg@20 ndcg@1 ndcg@10 ndcg@20
@@
w/ Role prompting 0.2750 0.4821 0.5379 0.2800 0.4814 0.5489
w/o CoT (step-by-step prompting) 0.2400 0.4543 0.5193 0.2650 0.4439 0.5163
w/ CoT (least-to-most prompting) 0.2783 0.5099 0.5603 0.2617 0.4729 0.5403
@@
w/ ICL (self) 0.2000 0.4558 0.5171 0.2500 0.4265 0.5055
w/ ICL (others) 0.1900 0.4015 0.4684 0.2500 0.4039 0.4950
@@ 5.4.3. Observations and Discussion
As shown in Table 12, findings on prompting strategies of LLMs are listed as follows:
@@
(1) The impacts on different expressions of prompts (RQ1). We compare different prompting strategies for LLMs, and report results based on the classification of zero-shot prompting and few-shot prompting. As shown in Table 12, “w/o” denotes that we remove related descriptions compared to the original prompt, and “w/” denotes that we add corresponding instructions to the prompt. As for the zero-shot prompting strategies, we can see that removing the recency-focused prompting sentence (w/o recency-focused) largely decreases the recommendation performance, demonstrating the key role of recently interacted items in recommendation. Although we provide historical items in chronological order based on timestamps, LLMs still needs explicit guidance to understand the importance of recent items, indicating that heuristic knowledge in the recommendation field needs to be supplemented for LLMs. Meanwhile, adding the role prompting descriptions to the original prompt (w/ role prompting) significantly improves the performance of zero-shot prompting, which shows that role-playing and expert-like prompts can better leverage the capabilities of LLMs in specific fields or tasks (Zhao et al., 2023). In addition, an important strategy for zero-shot prompting is the chain of thoughts, and we compare effects of two kinds of CoT prompting. When we remove the basic prompting sentence “let’s think step by step” in the original prompt, the recommendation effect is improved, possibly due to the fact that following the step-by-step prompts is not conducive to the extraction of results from textual outputs. It also indicates that specific problem decomposition may be required for recommendation tasks rather than general prompts, and the superior results of summarizing recent interest before recommendation (least-to-most prompting strategy) confirm this finding.
(2) The effects on planning modules of prompts (RQ2). In contrast to zero-shot prompting, the typical representative of few-shot prompting in ProLLM4Rec is in-context learning (ICL) based on contextual examples, and we study the ICL results considering demonstrations from the current user (self) and other users (others), respectively. From the results in Table 12, it can be seen that using examples of the target user is better than demonstrations from other users, indicating the personalized needs of each user. However, the strategy of few-shot prompting has insignificant advantage compared to the zero-shot prompting, and ICL is not fully applicable to recommendation scenarios.
@@
Observations of Prompting Strategies In chronological order, LLMs still needs explicit guidance to understand the importance of recent items. Role-playing and expert-like prompts can better leverage the capabilities of LLMs in specific fields. In chain-of-thought prompting, specific problem decomposition is required for recommendation tasks. The few-shot prompting strategy has insignificant advantages in recommendation scenarios.
@@ 6. Conclusion
This paper aims to provide a comprehensive exploration of Large Language Models (LLMs) to serve as recommender systems. It presents a systematic review of the advancements made in LLM-based recommendations, generalizing related work into multiple scenarios and tasks in terms of LLMs and prompts. We also conduct extensive experiments on two public datasets to investigate empirical findings for recommendation with LLMs. Our objective is to assist researchers in gaining a deeper understanding of the characteristics, strengths, and limitations of LLMs utilized as recommender systems. Considering the significant progress in LLMs, the development of LLM-based recommendations holds the potential to better align the powerful capabilities of LLMs with the evolving needs of intended users in the field of recommender systems. By addressing current challenges, we hope that our work will contribute to the advancement of LLM-based recommendations and serve as an inspiration for future research efforts. Last but not least, we outline promising directions for future research in utilizing LLMs for recommendation as follows.
@@
$\bullet$ Efficiency optimization of LLMs for recommendation. The key limitation of leveraging LLMs in industrial recommender systems is efficiency(Li et al., 2023l; Wu et al., 2023b; Fan et al., 2023), including considerations of both time and space. On the one hand, the fine-tuning and inference efficiency of LLMs cannot compare to traditional recommendation models (Zheng et al., 2023; Hou et al., 2023). While techniques such as parameter-efficient fine-tuning can aid in keeping LLMs updated in a computationally efficient manner, recommender systems need to iterate continuously over time, i.e., incremental learning. Frequent updates of LLMs inevitably bring spatial and temporal burdens to recommender systems (Shi et al., 2023). On the other hand, billions of parameters in LLMs also pose challenges for the lightweight deployment of recommendation algorithms (Touvron et al., 2023a, b). Therefore, efficiency optimization of LLMs utilized as recommender systems is one of the prerequisites for large-scale applications, which has widespread application prospects and scientific research values (Shi et al., 2023; Wang et al., 2023b; Huang et al., 2023).
@@
$\bullet$ Knowledge distillation of LLMs for recommendation. Since LLMs as recommenders are limited by efficiency, another feasible approach is to distill (Sun et al., 2024; Li et al., 2023a) the recommendation capabilities of LLMs into lightweight models, striking a balance between efficiency and effectiveness. Specifically, knowledge distillation is a classic model compression method adopted in recommender systems (Sun et al., 2024; Tian et al., 2023), with the core idea of guiding lightweight student models to “imitate” teacher models with better performance and more complex structures such as LLMs. In the field of recommender systems, the collaborative optimization of LLMs and recommendation models can also be seen as the distillation process to inject knowledge from LLMs to traditional recommenders, enhancing representations of users and items from semantic features (Li et al., 2023a; Qiu et al., 2023; Xi et al., 2023). Due to the fact that knowledge distillation can improve efficiency while retaining the capabilities of LLMs for recommendation, more applications need to be fully explored.
@@
$\bullet$ Multimodal recommendations with LLMs. In addition to IDs and text, multimodal recommendations with LLMs hold considerable promise and warrant comprehensive exploration with the evolving landscape of media consumption (Yang et al., 2023c; Zhang et al., 2023g). The essence of multimodal recommendations resides in the fusion of textual and visual information for enhanced user engagement (Pan et al., 2022; Harrison et al., 2023), and the dual functionality of LLMs is pivotal in this context. LLMs possess the capability to function as multimodal LLMs, enabling the incorporation and encoding of visual information extracted from images. Furthermore, images can be transformed into textual representations by multimodal encoders first, and LLMs are mainly used for the subsequent integration of diverse modalities (Harrison et al., 2023). In addition, the rich multimodal attributes also provide a basis for diversified recommendation results (Lin et al., 2023c). As the field progresses, an emphasis on the reproducibility, benchmarking, and standardization of evaluation datasets and metrics will be essential to foster a cohesive and informed advancement in multimodal recommender systems leveraging LLMs.
@@
$\bullet$ Fairness-aware recommendations of LLMs. Future research in the domain of fairness-aware recommendations with LLMs presents a compelling avenue for scholarly inquiry (Wang et al., 2023d; Zhang et al., 2023a; Li et al., 2023k; Dai et al., 2023b). In line with existing work, researchers have explored that retrievers in information retrieval are biased towards contents generated by LLMs (Dai et al., 2023b), and LLMs utilized as recommender systems also output unfair recommendation results (Wang et al., 2023d; Zhang et al., 2023a; Li et al., 2023k), necessitating a profound investigation into methodologies that ensure equitable and unbiased outcomes for LLM-based recommendations. It is imperative to scrutinize existing fairness-aware algorithms and develop novel approaches that cater to the intricacies of language models, particularly in understanding how fairness metrics align with user expectations. As LLMs continue to evolve, the research community must collaborate to establish standardized evaluation metrics and benchmarks for fairness-aware recommendations, ensuring the reproducibility and comparability of findings across diverse studies in the era of LLMs. In essence, fairness-aware recommendations with LLMs are poised to contribute substantially to the development of ethical and equitable recommender systems (Wang et al., 2023d; Zhang et al., 2023a).
@@
$\bullet$ General-purpose LLMs in the vertical field of recommender systems. Developing a comprehensive framework for LLMs to address multiple recommendation tasks represents a significant avenue for scholarly exploration (Geng et al., 2022; Chu et al., 2023; Zhang et al., 2023f; Wang et al., 2023c). The endeavor involves formulating a unified and versatile structure that accommodates the intricacies of various recommendation tasks (Dai et al., 2023a), encompassing diverse modalities and user preferences (Lin et al., 2023c; Harrison et al., 2023). In addition, using agents of LLMs to simulate recommendation scenarios and make dynamic decisions themselves is also a feasible application of general-purpose LLMs (Wang et al., 2023i; Zhang et al., 2023c; Wang et al., 2023g). Research efforts should focus on refining the architecture, training methodologies, and adaptability of such a general framework to ensure optimal performance across different recommendation domains. Investigating transfer learning techniques within this framework, enabling the transfer of knowledge between recommendation tasks, is crucial for enhancing efficiency and leveraging shared information (Zhang et al., 2023f; Bao et al., 2023b). In general, general-purpose LLMs for recommendation have the potential to revolutionize recommender systems by providing a unified and scalable solution capable of addressing the multifaceted challenges posed by diverse recommendation tasks.
@@
$\bullet$ Privacy and ethical concerns. Prior studies (Shen et al., 2023; Weidinger et al., 2021) have highlighted the potential issue of language models generating unreliable or personal contents based on certain prompts and insecure instructions. Recommendation systems involve massive amounts of the user data (Ni et al., 2019; Guo et al., 2017), and it is crucial to remove private and potentially harmful information stored in LLMs to enhance the privacy and security of LLM-based applications (Carranza et al., 2023; Lei et al., 2023). Notably, researchers have observed that Reinforcement Learning from Human Feedback (RLHF) and model editing techniques (Geva et al., 2022) have the potential to restrain the generation of poisonous or harmful contents from LLMs, thereby mitigating privacy and ethical concerns associated with privacy-preserving recommender systems (Carranza et al., 2023). Nevertheless, it is imperative to acknowledge that the alignment technology of LLMs may be vulnerable to misuse. The apprehension exists that LLMs could be manipulated by malicious users to selectively influence agents and amplify specific viewpoints within recommender systems. Consequently, addressing the security and privacy implications of LLM-based recommendations is crucial, and there is a need to formulate public regulations to ensure responsible usage and mitigate potential risks (Weidinger et al., 2021).
Acknowledgements

### v2

@@ 5.4. Prompting Strategies
With the construction of prompts for recommendation tasks, prompting strategies are leveraged to further elicit the general abilities of LLMs in text comprehending and problem solving (Zhao et al., 2023; Liu et al., 2023e; Scao and Rush, 2021). In this section, we first summarize prompting strategies specialized for recommender systems, and provide empirical findings on two research questions.
@@ 5.4.1. Prompting Strategies Specialized for Recommender Systems
The planning ability emerging from LLMs is the key to distinguishing them from other language models (Zhao et al., 2023). In general tasks with LLMs, researchers use methods such as Chain-of-Thought (CoT) prompting to stimulate the logical reasoning ability of LLMs for solving complex problems (Diao et al., 2024; Liu et al., 2022). As for LLM-RS, leveraging LLMs for recommender systems can also employ prompting strategies to further improve the performance of recommendations (Yao et al., 2023; Liu et al., 2024d). In addition, due to the user and item settings of the recommendation task, the prompting strategy needs to be customized based on the specific user needs in recommender systems. In this paper, we concentrate on typical prompting strategies for recommendation tasks, i.e., zero-shot prompting, few-shot prompting, recency-focused prompting, role prompting, chain-of-thought prompting and self-prompting strategy, and more advanced planning strategies in LLM-RS are left for future exploration (Liu et al., 2023e).
@@
$\bullet$ Zero-shot prompting is the basic scenario and common form among various prompting strategies. Zero-shot prompting directly provides the context information and task description for LLMs without reference examples. The task description and prompt formats vary in line with different recommendation tasks (Wang et al., 2024c; Liu et al., 2023c; Dai et al., 2023a).
@@
$\bullet$ Few-shot prompting is opposite to zero-shot prompting, and it provides a few demonstrations in prompts to help LLMs better understand the user intention (Brown et al., 2020). While for recommendation tasks, the matching of provided examples with the current interest of personalized users determines the effectiveness of few-shot prompting. Therefore, it is worth discussing how to provide reliable context for subsequent recommendations (Hou et al., 2024).
@@
$\bullet$ Recency-focused prompting is first proposed by (Hou et al., 2024) based on the fact that the next predicted item in recommendations has a greater correlation with the recently interacted item than other historical items. Therefore, explicitly emphasizing the recently interacted items in prompts is also a practical prompting strategy. In our initial prompt, recency-focused prompting is used by default.
@@
$\bullet$ Role prompting refers to playing a “role-playing” game with the language model. In the recommendation task, we can specify the role of LLMs as the recommender system to serve users in a targeted manner, adding auxiliary information to describe the role of recommenders in detailed prompts such as conversational recommender systems (He et al., 2023; Wang et al., 2023f; Jin et al., 2023).
@@
$\bullet$ Chain-of-thought prompting is also known as the CoT prompting, which is widely applied in reasoning tasks such as question answering and mathematical inference (Wei et al., 2022). CoT prompting can elicit the ability of LLMs to solve problems step by step (Kojima et al., 2022), or further explicitly decompose the reasoning and analysis process of the task using least-to-most prompting strategy (Zhou et al., 2023). In the field of recommender systems, explicit steps can also be provided manually by researchers to assist in solving recommendation tasks (Wang and Lim, 2023). For LLM-RS, we not only focus on the basic CoT prompts, but also explore the role of custom-designed step-by-step prompts.
@@
$\bullet$ Self-prompting strategy means that the knowledge for answering questions can be obtained by prompting LLMs multiple times. LLMs can be required to generate relevant knowledge, providing necessary information for concepts in the original problem (Liu et al., 2022). Meanwhile, the randomness and self consistency (Wang et al., 2023g) generated by LLMs allow them to generate multiple inference chains, and then use the majority voting method on the results obtained from all chains as final predictions. When using LLMs to the re-ranking stage, there are many possible combinations of candidate items, and the recommendation results from the same LLM with the same prompt can be totally different. Therefore, bootstrapping with multiple trials is a scientific guarantee to reduce bias (Hou et al., 2024; Ma et al., 2023; Li et al., 2023h).
@@ 5.4.2. Research Questions and Experimental Setup
In this section, we explore the construction and design of prompting strategies for LLM-RSon the following two research questions:
  • RQ1: When prompting LLMs a zero-shot recommender systems, what are the performance differences of different prompting strategies?
  • RQ2: When fine-tuning LLMs as recommender systems, what are the performance differences of different prompting strategies?
Figure 11. Examples of prompting strategies for LLMs in the scenario of the sequential re-ranking task based on Fig. 2(a). “ICL” denotes In-Context Learning and “CoT” stands for Chain-of-Thoughts.
[…]
Examples of prompting strategies for LLMs in the scenario of the sequential re-ranking task based on 2(a). “ICL” denotes In-Context Learning and “CoT” stands for Chain-of-Thought.
In the scenario of the sequential re-ranking task based on Fig. 2(a), we conduct experiments with prompting strategies illustrated in Fig. 11, and the LLM is ChatGPT:
@@
  • w/o CoT (step-by-step): we remove the magical spell “let’s think step by step” in the original prompt to observe the performance differences.
@@
  • w/ ICL (self instruction): we use the last interaction of the same user as the example for ICL (Hou et al., 2024).
  • w/ ICL (demonstration from others): we retrieve an example from historical sequences of other users for the current user, and add the demonstration to the original prompt. For the selection of examples, we encode recent items of users by textual representations, and search for similar user sequences based on the inner product between vectors.
In the scenario of fine-Tuning LLMs as the CTR predictor based on Fig. 2(b), we employ LLaMA3-8B as the LLM and LoRA is utilized for PEFT. As for prompting strategies of fine-tuning LLaMA3-8B, we devise four prompts as follows:
  • Implicit prompting: it is the basic prompt used in our experiments as illustrated in Fig. 2(b). According to the rating threshold, we divide the items users interact with into likes and dislikes, and only provide the implicit feedback for LLMs as prompts.
  • Explicit prompting: compared to implicit prompting, explicit prompting does not divide items, and provides LLMs with explicit feedback of ratings scores. An example is provided in the third row of Table 14.
  • Hybrid implicit and explicit prompting: based on implicit prompting, hybrid prompting further adds ratings of the target user for each provided item. An example is provided in the fourth row of Table 14.
  • CoT prompting: we add the magical spell “let’s think step by step” to the basic prompt of implicit prompting.
Table 13. Performance comparison on prompting strategies in Fig. 11 when leveraging ChatGPT as re-ranker. “w/o” denotes that we remove related descriptions compared to the original prompt, and “w/” denotes that we add corresponding instructions to the prompt.
MovieLens-1M Amazon Books
Method N@1 N@10 N@20 N@1 N@10 N@20
@@
[…]
@@
w/ ICL (self instruction) 0.2000 0.4558 0.5171 0.2500 0.4265 0.5055
w/ ICL (demonstration from others) 0.1900 0.4015 0.4684 0.2500 0.4039 0.4950
Table 14. Performance comparison of prompting strategies for fine-Tuning LLaMA3-8B on the CTR prediction task. The model is fine-tuned using LoRA, and the recommendation performance is evaluated using the accuracy metric on the MovieLens-1M (Movie) and Amazon Books (Book) datasets.
Prompting Strategy Movie Book
Implicit Prompting (Given the user’s preference and unpreference, identify whether the user will like the target item by answering ”Yes.” or ”No.”.) 0.6577 0.8610
Explicit Prompting (Given the user’s ratings of items they have interacted with (1-5, with 1 being the lowest and 5 being the highest), identify $\ldots$) 0.6847 0.8990
Hybrid Implicit and Explicit Prompting (Given the user’s ratings of items they have interacted with, we consider ratings greater than or equal to 4 as preferences, and ratings less than 4 as unpreferences, identify $\ldots$) 0.6649 0.8978
[…]
@@ 5.4.3. Observations and Discussion
Findings on prompting strategies of LLMs are listed as follows:
@@
(1) The impact of prompting strategies on prompting LLMs (RQ1). As shown in Table 13, we compare different prompting strategies for ChatGPT as the zero-shot re-ranker, and report results based on the classification of zero-shot prompting and few-shot prompting. As for the zero-shot prompting strategies, we can see that removing the recency-focused prompting sentence (w/o recency-focused) largely decreases the recommendation performance, demonstrating the key role of recently interacted items in recommendation. Although we provide historical items in chronological order based on timestamps, LLMs still needs explicit guidance to understand the importance of recent items, indicating that heuristic knowledge in the recommendation field needs to be supplemented for LLMs. Meanwhile, adding the role prompting descriptions to the original prompt (w/ role prompting) significantly improves the performance of zero-shot prompting, which shows that role-playing and expert-like prompts can better leverage the capabilities of LLMs in specific fields or tasks (Zhao et al., 2023). In addition, an important strategy for zero-shot prompting is the CoT, and we compare effects of two kinds of CoT prompting. When we remove the basic prompting sentence “let’s think step by step” in the original prompt, the recommendation effect is improved, possibly due to the fact that following the step-by-step prompts is not conducive to the extraction of results from textual outputs. It also indicates that specific problem decomposition may be required for recommendation tasks rather than general prompts, and the superior results of summarizing recent interest (least-to-most prompting strategy) confirm this finding. In contrast to zero-shot prompting, the typical representative of few-shot prompting is ICL based on contextual examples, and we study the ICL results considering demonstrations from the current user (self) and other users (others), respectively. It can be seen that using examples of the target user is better than demonstrations from other users, indicating the personalized needs of each user. However, the strategy of few-shot prompting has insignificant advantage compared to the zero-shot prompting, and more advanced planning strategies such as automatic prompting (Li et al., 2024a) and soft prompting (Ramos et al., 2024) shed lights on LLM-based recommender systems.
(2) The impact of prompting strategies on fine-tuning LLMs (RQ1). As shown in Table 14, we compare four prompting strategies for fine-Tuning LLaMA3-8B with LoRA on the CTR prediction task, highlighting the importance of prompt design in optimizing LLMs. Explicit prompting is the most effective strategy since it provides LLMs with quantifiable user preferences, enabling LLMs to capture nuanced user-item interactions. In contrast, Implicit prompting that relies on two-category data of user preferences gets poor results, suggesting that the lack of explicit guidance may hinder the capabilities of LLMs to infer user interest accurately. The hybrid implicit and explicit prompting strategy, while combining elements of both approaches, does not surpass explicit prompting, indicating that it is not scientific to set the threshold of likes and dislikes artificially. CoT prompting with step-by-step reasoning demonstrates competitive performance, suggesting that incorporating reasoning pathways can enhance the decision-making process when fine-tuning LLMs. These results emphasize that the choice of prompting strategy is not merely a technical detail but a critical factor influencing the effectiveness of fine-tuning LLMs. Future work could explore more sophisticated hybrid approaches or adaptive prompting techniques that dynamically adjust based on the complexity of the input or the task at hand. Additionally, investigating the interplay between prompting strategies and model architectures could provide deeper insights into optimizing LLMs for recommendation tasks (Tan et al., 2024; Yin et al., 2024).
@@
Observations of Prompting Strategies Although provided with historical items in chronological order, LLMs still needs explicit guidance to understand the importance of recent items. Role-playing and expert-like prompts can better leverage the capabilities of LLMs. In chain-of-thought prompting, specific problem decomposition is required for recommendation tasks rather than general prompts. The few-shot prompting strategy has insignificant advantages in recommendations. When fine-tuning LLMs as recommender systems, the explicit feedback (e.g., ratings) is better than the implicit feedback (e.g., interactions) to elicit capacities of LLMs.
@@ 6. Conclusion
This paper aims to provide a comprehensive exploration of Large Language Models (LLMs) to serve as recommender systems. It presents a systematic review of the advancements made in LLM-based recommendations, generalizing related work into multiple scenarios and tasks in terms of LLMs and prompts. We also conduct extensive experiments on two public datasets to investigate empirical findings for recommendation with LLMs. Our objective is to assist researchers in gaining a deeper understanding of the characteristics, strengths, and limitations of LLMs leveraged as recommender systems. Considering the significant progress in LLMs, the development of LLM-based recommendations holds the potential to better align the powerful capabilities of LLMs with the evolving needs of intended users in the field of recommender systems. By addressing current challenges, we hope that our work will contribute to the advancement of LLM-based recommendations and serve as an inspiration for future research efforts. Last but not least, we outline promising directions for future research in leveraging LLMs for recommendation.
@@
$\bullet$ Efficiency optimization of LLMs for recommendation. The key limitation of leveraging LLMs in industrial recommender systems is efficiency(Li et al., 2024d; Wu et al., 2024b; Zhao et al., 2024), including considerations of both time and space. On the one hand, the fine-tuning and inference efficiency of LLMs cannot compare to traditional recommendation models (Zheng et al., 2024; Hou et al., 2024). While techniques such as parameter-efficient fine-tuning can aid in keeping LLMs updated in a computationally efficient manner, recommender systems need to iterate continuously over time, i.e., incremental learning. Frequent updates of LLMs inevitably bring spatial and temporal burdens to recommender systems (Shi et al., 2024). On the other hand, billions of parameters in LLMs also pose challenges for the lightweight deployment of recommendation algorithms (Touvron et al., 2023a, b). Therefore, efficiency optimization of LLMs utilized as recommender systems is one of the prerequisites for large-scale applications, which has widespread application prospects and scientific research values (Shi et al., 2024; Wang et al., 2023b; Huang et al., 2023).
@@
$\bullet$ Knowledge distillation of LLMs for recommendation. Since LLMs as recommenders are limited by efficiency, another feasible approach is to distill (Sun et al., 2024; Li et al., 2023a) the recommendation capabilities of LLMs into lightweight models, striking a balance between efficiency and effectiveness. Specifically, knowledge distillation is a classic compression method adopted in recommender systems (Sun et al., 2024; Tian et al., 2023), with the core idea of guiding lightweight student models to “imitate” teacher models with better performance and more complex structures such as LLMs. In recommender systems, the collaborative optimization of LLMs and recommendation models can also be seen as the distillation process to inject knowledge from LLMs to traditional recommenders, enhancing representations of both users and items (Li et al., 2023a; Qiu et al., 2023; Xi et al., 2023). Due to the fact that knowledge distillation can improve efficiency while retaining the capabilities of LLMs for recommendation, more applications need to be fully explored.
@@
$\bullet$ Multimodal recommendations with LLMs. In addition to IDs and text, multimodal recommendations with LLMs hold considerable promise and warrant comprehensive exploration with the evolving landscape of media consumption (Yang et al., 2023c; Zhang et al., 2024c). The essence of multimodal recommendations resides in the fusion of textual and visual information for enhanced user engagement (Pan et al., 2022; Harrison et al., 2023), and the dual functionality of LLMs is pivotal in this context. LLMs possess the capability to function as multimodal LLMs, enabling the incorporation and encoding of visual information extracted from images. Furthermore, images can be transformed into textual representations by multimodal encoders first, and LLMs are mainly used for the subsequent integration of diverse modalities (Harrison et al., 2023). In addition, the rich multimodal attributes also provide a basis for diversified recommendation results (Lin et al., 2023b). As the field progresses, an emphasis on the reproducibility, benchmarking, and standardization of evaluation datasets and metrics will be essential to foster a cohesive and informed advancement in multimodal recommender systems leveraging LLMs (Geng et al., 2023; Wei et al., 2024a).
@@
$\bullet$ Fairness-aware recommendations of LLMs. Future research in the domain of fairness-aware recommendations with LLMs presents a compelling avenue for scholarly inquiry (Wang et al., 2024d; Zhang et al., 2023a; Li et al., 2023j; Dai et al., 2023b). In line with existing work, researchers have explored that retrievers in information retrieval are biased towards contents generated by LLMs (Dai et al., 2023b), and LLMs utilized as recommender systems also output unfair recommendation results (Wang et al., 2024d; Zhang et al., 2023a; Li et al., 2023j), necessitating a profound investigation into methodologies that ensure equitable and unbiased outcomes for LLM-based recommendations. It is imperative to scrutinize existing fairness-aware algorithms and develop novel approaches that cater to the intricacies of language models, particularly in understanding how fairness metrics align with user expectations. As LLMs continue to evolve, the research community must collaborate to establish standardized evaluation metrics and benchmarks for fairness-aware recommendations, ensuring the reproducibility and comparability of findings across diverse studies in the era of LLMs. In essence, fairness-aware recommendations with LLMs are poised to contribute substantially to the development of ethical and equitable recommender systems (Wang et al., 2024d; Zhang et al., 2023a).
@@
$\bullet$ General-purpose LLMs in the vertical field of recommender systems. Developing a comprehensive framework for LLMs to address multiple recommendation tasks represents a significant avenue for scholarly exploration (Geng et al., 2022; Chu et al., 2023; Zhang et al., 2023e; Wang et al., 2024c). The endeavor involves formulating a unified and versatile structure that accommodates the intricacies of various recommendation tasks (Dai et al., 2023a), encompassing diverse modalities and user preferences (Lin et al., 2023b; Harrison et al., 2023). In addition, using agents of LLMs to simulate recommendation scenarios and make dynamic decisions is also a feasible application of general-purpose LLMs (Wang et al., 2023h; Zhang et al., 2024a; Wang et al., 2023e). Research efforts should focus on refining the architecture, training methodologies, and adaptability of such a general framework to ensure optimal performance across different recommendation domains. Investigating transfer learning techniques within this framework, enabling the transfer of knowledge between recommendation tasks, is crucial for enhancing efficiency and leveraging shared information (Zhang et al., 2023e; Bao et al., 2023b). In general, general-purpose LLMs for recommendation have the potential to revolutionize recommender systems by providing a unified and scalable solution capable of addressing the multifaceted challenges.
@@
$\bullet$ Privacy and ethical concerns. Prior studies (Shen et al., 2023; Weidinger et al., 2021) have highlighted the potential issue of language models generating unreliable or personal contents based on certain prompts and insecure instructions. Recommendation systems involve massive amounts of the user data (Ni et al., 2019; Guo et al., 2017), and it is crucial to remove private and potentially harmful information stored in LLMs to enhance the privacy and security of LLM-based applications (Carranza et al., 2023; Lei et al., 2023). Notably, researchers have observed that Reinforcement Learning from Human Feedback (RLHF) and model editing techniques (Geva et al., 2022) have the potential to restrain the generation of poisonous or harmful contents from LLMs, thereby mitigating privacy and ethical concerns associated with privacy-preserving recommender systems (Carranza et al., 2023). Nevertheless, it is imperative to acknowledge that the alignment technology of LLMs may be vulnerable to misuse. The apprehension exists that LLMs could be manipulated by malicious users to selectively influence agents and amplify specific viewpoints within recommender systems. Consequently, addressing the security and privacy implications of LLM-based recommendations is crucial, and there is a need to formulate public regulations to mitigate potential risks (Weidinger et al., 2021).
Acknowledgements.